# PaintScene4D: Consistent 4D Scene Generation from Text Prompts

### v1

@@ D Limitations and Future Work
While our method successfully generates photorealistic 4D scenes from a single text prompt, several limitations persist. First, our approach assumes that the input video is captured from a nearly static, non-moving camera. This assumption does not always hold when using text-to-video (T2V) models, which typically offer limited control over camera dynamics. Videos with large camera movement results in a degraded visual quality. We demonstrate the failure case in Figure A. Therefore, extending our framework to accommodate videos with more camera movements represents a promising direction for future work. Second, our current method does not explicitly model the 3D structure of the foreground. Instead, we rely on an inpainting model to fill in gaps at the boundaries of the foreground, which means that the model does not possess a comprehensive understanding of the 3D geometry of the scene. A more advanced approach could involve explicitly separating the foreground from the background and modeling the 3D structure of the foreground, potentially using methods like SV4D [52]. Finally, our approach struggles to handle rapid movements in the video due to the limitations of current 4D rendering techniques. Advancements in this area would likely enhance the rendering quality of our method and enable better handling of fast motion.

### v2

@@ F Limitations and Future Work
    Assumption of a Static Camera: Our approach assumes that the input video is captured from a nearly static, non-moving camera. This assumption does not always hold when using text-to-video (T2V) models, which typically offer limited control over camera dynamics. Videos with a large camera movement result in a degraded visual quality. We demonstrate the failure case in Figure B. Therefore, extending our framework to accommodate videos with more camera movements represents a promising direction for future work.
  2. 2.
    Lack of Explicit 3D Foreground Modeling: Our current method does not explicitly model the 3D structure of the foreground. Instead, we rely on an inpainting model to fill in gaps at the boundaries of the foreground, which means that the model does not possess a comprehensive understanding of the 3D geometry of the scene. A more advanced approach could involve explicitly separating the foreground from the background and modeling the 3D structure of the foreground, potentially using methods like SV4D [53].
  3. 3.
    Challenges with Rapid Motion: Our approach struggles to handle rapid movements in the video due to the limitations of current 4D rendering techniques. Advancements in this area would likely enhance the rendering quality of our method and enable better handling of fast motion.
  4. 4.
    Segmentation Errors and Artifacts: If the segmentation model fails to accurately distinguish the character or foreground from the background, it can introduce significant errors during the warping and inpainting processes. These inaccuracies accumulate over successive stages, leading to noticeable artifacts in the final rendered output. One common issue is the presence of double geometry, where duplicated or misaligned structures appear in the scene, reducing the overall visual quality and realism of the generated 4D representation.
Figure C: Gallery of Results: We present qualitative results of our text-to-4D generation framework, showcasing superior visual fidelity, consistent multi-view reconstructions, plausible scene compositions, and realistic dynamic motions. The horizontal axis represents the time axis and the vertical axis represents different viewpoints. A comprehensive collection of video demonstrations is provided in the supplementary materials.
Figure D: Gallery of More Results. A comprehensive collection of video demonstrations is provided in the supplementary materials.